# Augmenting Smart Contract Decompiler Output through Fine-grained Dependency Analysis and LLM-facilitated Semantic Recovery

### v1

@@ 3. Design of SmartHalo
Figure 3. An optimization error reported by LLM inference in terms of program-behavior non-equivalence.
@@
To overcome the deficiencies, we propose a novel decompiler output optimization framework, called SmartHalo, based on hybrid SA and LLM. SmartHalo takes the decompiler output as input, and finally outputs the optimized code for analysts. Fig. 4 shows the overview of SmartHalo, which integrates the following mechanisms to ensure its effectiveness.
@@
Dependency-based Semantic Extraction.  In this mechanism, SmartHalo aims to extract crucial semantics from the decompiled code through a fine-grain dependency analysis, which is then propagated into the LLM for further processing. To this end, our work complements and extends previous research by considering two new types of dependencies: namely, state dependency and type dependency. Specifically, SmartHalo analyzes three types of dependencies as the crucial program semantics, subsequently utilizing them to construct a Dependency Graph (DG). The extracted semantics (i.e., dependencies) represent the critical static domain knowledge necessary for decompiler output optimization, which facilitates the LLMs in capturing them across the entire smart contract (See Section 4.1).
@@
LLM-driven Semantic Enrichment.  With the semantic re-presentation (i.e., DG) propagated by the above mechanism, SmartHalo constructs inference prompts to guide LLM to enrich the semantic of decompiler output. Specifically, according to the DG, SmartHalo utilizes our designed generator to constructs the context for optimization targets (i.e., methods or variables), for determining the code range related to optimization targets. Subsequently, SmartHalo further generates chain-of-thoughts that represent inference steps of static analysis for optimization targets, based on the dependencies among the DG, for teaching LLMs how to infer variable types, contract attributes and method boundaries. In addition, SmartHalo collects the candidates for variable types and contract attributes from the Solidity documentation (Sol, [n. d.]). Lastly, LLM learns from the above three types of prompt, and optimizes the crucial semantic (i.e., variable types, contract attributes and method boundaries) for decompiled code (See Section 4.2).
Correctness Verification.  To maintain the fidelity of the optimized code, SmartHalo should ensure that the optimized code preserves the consistent program behaviors with the original decompiler output and eliminate any inference errors that breach static domain rules.
@@
However, verifying correctness is far from straightforward. First, due to the hallucination that commonly exist in LLMs, LLMs generates the random and unpredictable outputs that unexpectedly alter the program behaviors of the original decompiled code. For example, Fig. 3 illustrates a program behavior non-equivalence error caused by LLM optimization. In the original decompiled code, variables v0, v3 and v9 depends on parameters varg6, varg7 and varg8, respectively (line 02-04). However, in the optimized code, variables sigV, sigR and sigS incorrectly depend on the same parameter signaturesCount (line 06-08). Second, in addition to program behavior non-equivalence, LLMs also introduce inference errors that contradict common static domain rules, further complicating correctness verification. We take Fig. 3 again as an instance for illustrating such type of inference errors. In line 18, the value returned by the predefined function keccak256 is type of bytes32. According to Solidity documentation (Sol, [n. d.]), bytes32 cannot be directly converted into uint256, result in convertion errors.
@@
SmartHalo implements two specific checks for establishing correctness verification. In the program-behavior equivalence check, SmartHalo utilizes symbolic execution and formal verification to verify the program-behavior equivalence between the original decompiled code and optimized code. In the static rule violation check, SmartHalo summarizes a set of static rules that are related to types. Note that static rule violation check donot consider method boundaries and contract attributes because there are no explicit static rules related to them in smart contract (Sol, [n. d.]). By applying these static rules, SmartHalo rejects the inference errors. (see Section 4.3)
Figure 4. The overview of SmartHalo.
4. Approach Details
4.1. Dependency-based Semantic Extraction
@@
  • State dependency. State dependency refers to the dependency between different state variables or dependency between state variables and expressions, such as read and write dependency on state variables. SmartHalo utilize SmartState (Liao et al., 2023), a SOTA state dependency analyzer to identify state dependency from given smart contracts.
Contract Dependency Graph Construction.  Subsequently, SmartHalo proposes the dependency graph (DG) as the representation of static domain knowledge for smart contracts.
@@
Figure 5. The syntax of expression for typing in Solidity.
@@
Figure 6. The optimization prompt for the instance in Figure 1 (b).
4.2. LLM-driven Semantic Enrichment
@@
Code Context.  SmartHalo adapts different strategies to generate code snippets for variables and methods. The strategy for variables applies to optimizing variable type and contract attribute, and another strategy for methods is used to optimize method boundaries.
$\bullet$ Variables. To pinpoint the context related to target variables precisely, SmartHalo utilizes code slicing to identify code context fragments from the decompiled code for target variables, based on the DG. To this end, SmartHalo first generates the slicing DG based on the original DG. On the original DG, SmartHalo starts from target variable nodes and performs the forward and backward traversals (i.e., the same or opposite direction to DG edges). In this way, SmartHalo finds out all the nodes that have a dependency on target variables and generates a slicing subgraph of DG. With the slicing DG, SmartHalo combines the corresponding expressions together to generate a code snippet as the context prompt. In this way, SmartHalo narrows down the original decompiled code to the much smaller code snippet, which contains the information only related to decompiler output optimization of target variables.
$\bullet$ Methods. To generate the context prompt for target methods, SmartHalo searches the DG to find out the call chains that target methods lie in. Then, SmartHalo combines the corresponding methods in these call chains together to generate a code snippet as the context prompt.
We present the code context generation for the instance of Fig. 1(b), as highlighted in the gray part of Fig. 6. SmartHalo identifies all the expressions that have type dependency related to variable setStorage, and integrates them as the code context prompt.
@@
Variable type candidates. Solidity is a statically typed language that provides a specific number of built-in types. In addition, the user-defined value type in Solidity is actually an alias that creates a zero cost abstraction over an elementary value type, so SmartHalo directly utilizes the elementary value type to represent it. According to Solidity documentation (Sol, [n. d.]), we collect the variable type candidates, as can be seen in the inference candidates (highlighted in green part) of Fig. 6.
@@
Contract attribute candidates. In contrast to variable type, contract attribute is far from straightforward and should be summarized by representative smart contract dataset. We summarize the contract attribute candidates as a set [ Limit, Fee, Flag, Address, Asset, Router, Others ]  [^1] The details of manual summarizing process for contract attribute candidates can be assessed via Section Data Availability., and the detail of summarizing process are illustrated in Section 5.1
@@
Chain-of-Thought Prompt.  COT prompts are actually a series of intermediate reasoning steps in static analysis. SmartHalo prompts the LLM with COT Prompt, for teaching LLM how to reason the correct method boundaries, variable type, and contract attribute from the perspective of static domain knowledge.
@@
SmartHalo utilizes the slicing DG of the code slicing phase to generate the COT prompts. Given the slicing DG, SmartHalo organizes the nodes according to the distance between the nodes and the target variable node (i.e., hop counts). Specifically, the hop count of the target variable node is set as 0, and the hop counts of other nodes are calculated by their distance. SmartHalo translates each hop in the DG into a sentence of dependency description. Inspired by prior work (Peng et al., 2023), we summarize the effective COT prompt templates, as shown in Table 1. As can be seen, SmartHalo utilizes different COT prompt templates for the different optimization tasks. As can be seen, the category of type dependency is designed for the optimization on variable type recovery. Further, the category of state dependency is specific to the optimization on contract attribute recovery. The category of control flow dependency is used for the optimization on method boundary identification. To generate the COT prompt, SmartHalo fills the corresponding variable type, variable name, usages and statement, into the templates.
@@
To illustrate the COT prompt generation, we present the COT prompt for the instance of Fig. 1(b), as highlighted in yellow part of Fig. 6. As discussed above, there are three type dependency edges (i.e., $D_{t_{1}}$, $D_{t_{2}}$, $D_{t_{3}}$) for the instance of Fig. 1(b). Among them, SmartHalo utilizes $expression\rightarrow variable$ template to describe $D_{t_{1}}$, and leverages $variable\rightarrow variable$ template to describe $D_{t_{2}}$ and $D_{t_{3}}$. The sentences of all the corresponding edges are connected together according to the order, to form the COT prompt.
Table 1. Chain-of-Thought Prompt Template for LLM-driven semantic enrichment
Category Type Template
Type dependency Variable $\to$ Variable
The type of variable[NAME] is consistent with the type of variable [NAME]
Type $\to$ Expression
The value of predefined function/operands[NAME] of expression [STATEMENT] is type of [TYPE]
Type $\to$ Variable The variable of [NAME] is assigned from [TYPE]
Expression $\to$ Variable
The type of variable [NAME] depends on expression [STATEMENT]
Variable $\to$ Expression
The operand(s)/target(s)/key(s)/value(s) of expression [STATEMENT] is/are [TYPE].
State dependency State variable $\to$ State variable
The attribute of state variable [NAME] is correlated to the attribute of state variable [NAME]
State variable $\to$ Expression
State variable [NAME] is written by Expression [STATEMENT]
Expression $\to$ State variable
State variable [NAME] is read by Expression [STATEMENT]
Control flow dependency
Return Value
Expression [STATEMENT] is used for return the value
Variable declaration Expression [STATEMENT] is used to declare the variable
Variable Usages The variable [Name] is used for [USAGES]
[NAME] indicates the name of variable, [STATEMENT] indicates the STATEMENT of expression nodes,
[TYPE] indicates the name of type nodes. [USAGES] indicates the usages for the variable.
4.3. Correctness Verification

### v2

@@ II-B Deficiencies in Decompiler Output
Figure 3: An optimization error reported by LLM inference in terms of program-behavior non-equivalence.
II-C Our solution
We propose our strategies for optimizing the decompiler output, combining with a discussion on potential solutions for the deficiencies presented in motivating examples of Fig. 1.
Strategy-1: Using SA to extract control-flow dependency, type dependency and state dependency.  First, type dependency is crucial for the variable type recovery. For example, the type error for variable uintStorage can be corrected by using SA to infer type dependencies, as illustrated in Fig. 1(a). Initially, the output of the predefined function keccak256 is of type bytes32 ($D_{t_{1}}$). Subsequently, the type of variable $v0$ depends on the return value type of the function _getTokenKey, which is also bytes32 ($D_{t_{2}}$). Finally, the key type of the variable setStorage depends on the type of $v0$ ($D_{t_{3}})$. Consequently, the type of setStorage should be corrected to mapping(bytes32= uint256). Second, control-flow dependency is critical for the function boundary identification. For instance, as shown in Fig. 1(c), while static constraints (i.e., call information) are unavailable in function validate, we can manually infer its boundaries by analyzing the contextual control flow, i.e., the function typically declares a new variable at the beginning (line 10) and returns the value at the end (line 11). Third, state dependency is important for contract attribute inference. We take the motivating example in Fig. 1(b) as an instance for illustration. By analyzing the dependency of the state variable, we can find that state variable stor_0 is written in line 16 ($D_{s_{1}}$) and line 18 ($D_{s_{2}}$). By manually analyzing their context usage, we can infer that they are utilized for asset transfer and asset receive, respectively. Hence, we can infer the attribute for stor_0 as the [Asset].
Strategy-2: Empowering LLMs to enrich semantics.  With the remarkable capability of few-shot in-context learning [28, 29], LLMs are flexible to predicting the optimization target (i.e., variables or functions), which are rarely encountered in the program and often incapable of inferring them based on other variables or functions. We take Fig. 1(a) again as an instance for illustration. The type error for variable input can be corrected by using LLMs for prediction. Fig. 3 presents code optimized by LLM. Benefiting from the extensive training set and model capabilities, LLM can assign the variable input with string memory correctly (line 12). In addition, owing to its remarkable model capacity, LLMs are more effective and generalizable in learning static domain knowledge (i.e., dependencies) for optimization predictions, compared to SA’s manually crafting heuristic rules based on the same knowledge [30]. For instances, instead of manually summarizing heuristic rules, we can utilize LLMs to learning the control-flow context for function boundary prediction, and to analyze the contextual usage of state variables to infer their contract attributes.
Inspired by the above analysis, our objective is to harness the advantages of both SA and LLMs to optimize decompiler output. Therefore, SmartHalo extracts three types of program dependencies including type dependency, state dependency and control flow dependency within the program through SA. Subsequently, SmartHalo leverages the superior generalization capabilities of LLMs to optimize the decompiled code after learning these program dependencies.
Figure 4: The overview of SmartHalo.
III Design of SmartHalo
@@
To overcome the deficiencies, we propose a novel decompiler output optimization framework, called SmartHalo, based on hybrid SA and LLM. SmartHalo takes the decompiler output as input, and finally outputs the optimized code. Fig. 4 shows the overview of SmartHalo, which integrates the following mechanisms to ensure its effectiveness.
@@
Dependency-based Semantic Extraction.  In this mechanism, SmartHalo aims to extract critical semantics from the decompiled code through a fine-grain dependency analysis, which is then propagated into the LLM for further processing. To this end, our work complements and extends previous research [11, 12, 6, 5] by considering two new types of dependencies: namely, state dependency and type dependency. Specifically, SmartHalo analyzes three types of dependencies as the crucial program semantics, subsequently utilizing them to construct a Dependency Graph (DG). The extracted semantics (i.e., dependencies) represent the critical static domain knowledge necessary for decompiler output optimization, which facilitates the LLMs in capturing them across the entire smart contract (See Section IV-A).
@@
LLM-driven Semantic Enrichment.  With the semantic re-presentation (i.e., DG) propagated by the above mechanism, SmartHalo constructs inference prompts to guide LLM to enrich the semantic of decompiler output. Specifically, according to the DG, SmartHalo utilizes our designed generator to constructs the context for optimization targets (i.e., functions or variables), for determining the code range related to optimization targets. Subsequently, SmartHalo further generates chain-of-thoughts that represent inference steps of static analysis for optimization targets, based on the dependencies among the DG, for teaching LLMs how to infer variable types, contract attributes and function boundaries. In addition, SmartHalo collects the candidates for variable types and contract attributes from the Solidity documentation [3]. Lastly, LLM learns from the above three types of prompt, and optimizes variable types, contract attributes and function boundaries for decompiled code (See Section IV-B).
@@
Correctness Verification.  However, verifying correctness for the optimized code is far from straightforward. First, due to the hallucination that commonly exist in LLMs, LLMs generates the random and unpredictable outputs that unexpectedly alter the program behaviors of the original decompiled code. For example, Fig. 3 illustrates a program behavior non-equivalence error caused by LLM optimization. In the original decompiled code, variables v0, v3 and v9 depends on parameters varg6, varg7 and varg8, respectively (line 02-04). However, in the optimized code, variables sigV, sigR and sigS incorrectly depend on the same parameter signaturesCount (line 06-08). Second, LLMs also introduce errors that contradict common static domain rules (e.g., type rules), further complicating correctness verification. We take Fig. 3 again as an instance for illustrating such type of errors. In line 18, the value returned by the predefined function keccak256 is type of bytes32. According to Solidity documentation [3], bytes32 cannot be directly converted into uint256, result in convertion errors.
@@
SmartHalo implements two specific checks for establishing correctness verification. In the program-behavior equivalence check, SmartHalo utilizes symbolic execution and formal verification to verify the program-behavior equivalence between the original decompiled code and optimized code. In the static rule violation check, SmartHalo summarizes a set of static rules that are related to types. Note that static rule violation check donot consider function boundaries and contract attributes because there are no explicit static rules related to them in smart contract [3]. By applying these static rules, SmartHalo rejects the inference errors. (see Section IV-C)
IV Approach Details
IV-A Dependency-based Semantic Extraction
Basic control-flow analysis.  To enable control-flow analysis at the level of decompiled code, we develop a Solidity-oriented control-flow analysis component for SmartHalo, which provides both code parsing and control/data-flow graph construction. Built on a Tree-sitter architecture, the component performs analysis without requiring compilation, which eliminates environment configuration and build overhead. Such property makes the component particularly suitable for processing decompiled pseudocode. Specifically, the control-flow analysis component (i) employs Tree-sitter [31] to construct a concrete syntax tree for each decompiled code file, (ii) transforms the syntax tree into a three-address intermediate representation (IR), and (iii) leverages this IR together with function-call information to produce the corresponding control-flow and data-flow graphs.
@@ IV-A Dependency-based Semantic Extraction
  • State dependency. State dependency refers to the dependency between different state variables or dependency between state variables and expressions, such as read and write dependency on state variables. SmartHalo utilize SmartState [32], a SOTA state dependency analyzer to identify state dependency from given smart contracts.
Dependency Graph Construction.  Subsequently, SmartHalo proposes the dependency graph (DG) as the representation of critical semantics for smart contracts.
@@
Figure 5: The syntax of expression for typing in Solidity.
@@
IV-B LLM-driven Semantic Enrichment
@@
TABLE I: Chain-of-Thought Prompt Template for LLM-driven semantic enrichment
. Category Type Template Type dependency Variable $\to$ Variable The type of variable [NAME] is consistent with the type of variable [NAME] Type $\to$ Expression The value of predefined function/operands[NAME] of expression [STATEMENT] is type of [TYPE] Type $\to$ Variable The variable of [NAME] is assigned from [TYPE] Expression $\to$ Variable The type of variable [NAME] depends on expression [STATEMENT] Variable $\to$ Expression The operand(s)/target(s)/key(s)/value(s) of expression [STATEMENT] is/are [TYPE]. State dependency State variable $\to$ State variable The attribute of state variable [NAME] is correlated to the attribute of state variable [NAME] State variable $\to$ Expression The attribute of state variable [NAME] is correlated to the context usages of Expression [STATEMENT] (Write) Expression $\to$ State variable The attribute of state variable [NAME] is correlated to the context usages of Expression [STATEMENT] (Read) Control flow dependency Call Site Expression [STATEMENT] seems to be a call site to another function. Modifier Expressions with a range from [STATEMENT] to [STATEMENT] seems to be a modifier function. Return Value Expression [STATEMENT] is used to return the value. Please determine whether it is the end point. Variable Declaration Expression [STATEMENT] is used to declare the variable. Please determine whether it is the start point.
Note: [NAME] indicates the names of variable nodes, [STATEMENT] indicates the statements of expression nodes, [TYPE] indicates the types of variable nodes. [USAGES] indicates the usages for the variable nodes.
Code Context.  SmartHalo adapts different strategies to generate code snippets for variables and functions. The strategy for variables applies to optimizing variable type and contract attribute, and another strategy for functions is used to optimize function boundaries.
$\bullet$ Variables. To pinpoint the context related to target variables precisely, SmartHalo utilizes code slicing to identify code context fragments from the decompiled code for target variables, based on the DG. To this end, SmartHalo first generates the slicing DG based on the original DG. On the original DG, SmartHalo starts from target variable nodes and performs the forward and backward traversals (i.e., the same or opposite direction to DG edges). In this way, SmartHalo finds out all the nodes that have the dependency on target variables and generates a slicing subgraph of DG. With the slicing DG, SmartHalo combines the corresponding expressions together to generate a code snippet as the context prompt.
$\bullet$ functions. To generate the context prompt for target functions, SmartHalo searches the DG to find out the call chains that target functions lie in. Then, SmartHalo combines the corresponding functions in these call chains together to generate a code snippet as the context prompt.
In this way, SmartHalo narrows down the original decompiled code to the much smaller code snippet, which contains the information only related to decompiler output optimization of target variables or functions. We present the code context generation for the instance of Fig. 1, as highlighted in the gray part of Fig. 6. For type optimization, SmartHalo identifies all the expressions that have type dependency related to variable setStorage, and integrates them as the code context prompt. For attribute recovery, SmartHalo extracts the corresponding expressions that have state dependency on state variable stor_1, and aggregates them as the code context prompt. Finally, SmartHalo extract the functions from the call chain that function _handle\_handle lies in to form the code context prompt.
Figure 6: The prompt for the instance in Figure 1.
@@
Variable type candidates. Solidity is a statically typed language that provides a specific number of built-in types. In addition, the user-defined value type in Solidity is actually an alias that creates a zero cost abstraction over an elementary value type, so SmartHalo directly utilizes the elementary value type to represent it. According to Solidity documentation [3], we collect the variable type candidates, as can be seen in the inference candidates (highlighted in green part) of Fig. 6 (a).
@@
Contract attribute candidates. In contrast to variable type, contract attribute is far from straightforward and should be summarized by representative smart contract dataset. We summarize the contract attribute candidates as a set, as can be seen in the inference candidates (highlighted in green part) of Fig. 6 (b). And the detail of summarizing process are illustrated in Section V-A
$\frac{}{\pi\vdash c:\theta}$ (Constant)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad\pi\vdash e_{2}:\theta_{2},\quad\widetilde{\theta}=\left\{bool,int\right\}}{\pi\vdash e_{1}[bitop]e_{2}:\theta\wedge\widetilde{\theta},\quad\pi\vdash e_{1}:\theta_{1}\wedge\widetilde{\theta},\quad\pi\vdash e_{2}:\theta_{2}\wedge\widetilde{\theta}}$ (LShift, RShift)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad\widetilde{\theta}=\left\{bool,int\right\},\quad\pi\vdash e_{2}:\theta_{2},\quad{\theta}^{\prime}=getMorePreciseType(\theta_{1}\wedge\widetilde{\theta},\quad\theta_{2}\wedge\widetilde{\theta})}{\pi\vdash e_{1}[numop]e_{2}:{\theta}^{\prime},\quad\pi\vdash\theta_{1}\wedge\widetilde{\theta},\pi\vdash\theta_{2}\wedge\widetilde{\theta}}$ (Numeric Operations)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad\pi\vdash e_{2}:\theta_{2},\quad{\theta}^{\prime}=\left\{\Gamma,Array,Tuple\right\}}{\pi\vdash e_{1}[cmpop]e_{2}:bool,\quad\pi\vdash e_{1}:\theta_{1}\wedge\theta_{2}\wedge\widetilde{\theta},\quad\pi\vdash e_{2}:\theta_{1}\wedge\theta_{2}\wedge\widetilde{\theta}}$ (Lt, LtE, Gt, GtE)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad...,\quad\pi\vdash e_{n}:\theta_{n}}{\pi\vdash(e_{1},...,e_{n}):Tuple[\theta_{1},...,\theta_{n}],\quad\pi\vdash(e_{1},...,e_{n}):Array[\theta_{1},...,\theta_{n}]}$ (Tuple, Array)
$\frac{\pi\vdash e:\theta,\quad\widetilde{\theta}=\left\{str,bytes\right\},\quad{\theta}^{\prime}=getElementType(\theta_{1}\wedge\widetilde{\theta})}{\pi\vdash[for]v[in]e:{\theta}^{\prime},\quad\pi\vdash e:\theta\wedge\widetilde{\theta}}$ (Comprehension)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad\pi\vdash e_{2}:\theta_{2}}{\pi\vdash e_{1}[blop]e_{2}:{\theta}^{\prime},\quad Union[\theta_{1},\theta_{2}]}$ (Boolean Operation)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad\pi\vdash e_{2}:\theta_{2},\quad\widetilde{\theta}=\left\{bool,int,byte\right\}}{\pi\vdash e_{1}[bitop]e_{2}:\theta\wedge\widetilde{\theta},\quad\pi\vdash e_{1}:\theta_{1}\wedge\widetilde{\theta},\quad\pi\vdash e_{2}:\theta_{2}\wedge\widetilde{\theta}}$ (Bitor, BitAnd, BitXor)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad\pi\vdash e_{2}:\theta_{2}}{\pi\vdash e_{1}[cmpop]e_{2}:bool}$ (Eq,NotEq,Is,IsNot)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad\pi\vdash e_{2}:\theta_{2},...\quad\pi\vdash e_{n}:\theta_{n},\quad\widetilde{\theta}=\left\{Callable[[\theta_{1},\theta_{2},...,\theta_{n}],\theta]\right\},\quad{\theta}^{\prime}=getReturnType(\theta\wedge\widetilde{\theta})}{\pi\vdash e(e_{1},...,e_{n}):\theta}$ (Call)
$\frac{\pi\vdash e_{1}:\theta_{1},\quad\pi\vdash e_{2}:\theta_{2},\quad\widetilde{\theta_{1}}=\left\{str,bytes\right\},\quad\widetilde{\theta_{2}}=\left\{int,bool\right\},\quad{\theta}^{\prime}=getElementType(\theta_{1}\wedge\theta_{2})}{\pi\vdash e_{1}[e_{2}]:{\theta}^{\prime},\quad\pi\vdash e_{1}:\theta_{1}\wedge\widetilde{\theta_{1}},\quad\pi\vdash e_{2}:\theta_{2}\wedge\widetilde{\theta_{2}}}$ (Slice)
Figure 7: The static rules for type violation check.
@@
Chain-of-Thought Prompt.  The COT prompts are a series of intermediate reasoning steps in static analysis. SmartHalo prompts the LLM with COT Prompt, for teaching LLM how to reason the correct function boundaries, variable type and contract attribute from the perspective of static domain knowledge.
@@
SmartHalo utilizes the slicing DG of the code slicing phase to generate the COT prompts. Given the slicing DG, SmartHalo organizes the nodes according to the distance between the nodes and the target variable node (i.e., hop counts). Specifically, the hop count of the target variable node is set as 0, and the hop counts of other nodes are calculated by their distance. SmartHalo translates each hop in the DG into a sentence of dependency description. Inspired by prior work [30], we summarize the effective COT prompt templates, as shown in Table I. As can be seen, each optimization tasks involve at least one kind of COT prompt template. Specifically, variable type optimization relies on the categories of both type dependency and control-flow dependency. Further, variable type optimization depends on the categories of both static dependency and control-flow dependency. And function boundary optimization only requires the category of control-flow dependency. To generate the COT prompt, SmartHalo fills the corresponding variable [TYPE], variable [NAME], [USAGE] and [STATEMENT] into the templates.
@@
To illustrate the COT prompt generation, we present the COT prompt for the instance of Fig. 1, as highlighted in yellow part of Fig. 6. For type optimization, as discussed above, there are three type dependency edges (i.e., $D_{t_{1}}$, $D_{t_{2}}$, $D_{t_{3}}$) for the instance of Fig. 1(a). As shown in Fig. 6(a), SmartHalo utilizes $Expression\rightarrow Variable$ template to describe $D_{t_{1}}$, and leverages $Variable\rightarrow Variable$ template to describe $D_{t_{2}}$ and $D_{t_{3}}$. The sentences of all the corresponding edges are connected together according to the order of type dependency, to form the COT prompt for type optimization.
For contract attribute recovery, as mentioned above, there are two dependency edges (i.e., $D_{s_{1}}$, $D_{s_{2}}$) for the instance of Fig. 1(b). As shown in Fig. 6(b), SmartHalo utilizes $Statevariable\rightarrow Expression$ template to describe $D_{s_{1}}$ and and $D_{s_{2}}$. The sentences of these edges are aggregated together as the COT prompt for attribute recovery.
Fig. 6(c) shows the COT prompt for function boundary optimization on the instance of Fig. 1(c). SmartHalo utilizes ReturnValueReturn-Value and VariableDeclarationVariable-Declaration templates to describe statements that are likely to be the start and end points of a function.
IV-C Correctness Verification